OpenStudy (albert0898):

10. Line m is tangent to circle O at the point (12, -9.) If the center of O is at the origin, what is the y-intercept of line m? Note: I have already established that the radius is 15. Thank you :)

OpenStudy (anonymous):

What you can do in this situation is 

OpenStudy (anonymous):

Any tangent line is perpendicular to the radius of a circle So what is the slope of the radius in this case?

OpenStudy (albert0898):

@Brill The slope is -3/4

OpenStudy (anonymous):

ok so the slope of a line perpendicular to a line is the NEGATIVE RECIPROCAL So for example if you had a line of slope 20, the slope of a line perpendicular to that would be -1/20

OpenStudy (anonymous):

So what would the slope of the line be?

OpenStudy (albert0898):

4/3

OpenStudy (albert0898):

@Brill

OpenStudy (anonymous):

Yup, now you can input that into point slope form, input the point and find the y intercept

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (albert0898):

@Brill Well, point-slope form is y − y1 = m(x − x1)

OpenStudy (albert0898):

y+9 = 4x/3 - 48/3

OpenStudy (albert0898):

y = 4x/3 - 25

OpenStudy (anonymous):

Yup So you would get \[y-(-9)=\frac{ 4 }{ 3 }(x-12)\]

OpenStudy (albert0898):

Thank you for your help @Brill :)

OpenStudy (anonymous):

You want the y intercept, and at the y intercept x always equals 0 So sub 0 for x and solve for y

OpenStudy (anonymous):

No problem :)